# Model Restrictiveness in Functional and Structural Settings

### v2

@@ 5.3 Multinomial Choice with Endogeneity
Unlike in Section 5.2, we now allow price to be endogenous. We address this endogeneity using the BLP instruments constructed in Nevo (2000), selecting the two instruments that are most strongly correlated with price from the full set of 20.  [^8] In Online Appendix F.3, we also consider selecting the three instruments most strongly correlated with price; the qualitative ranking of restrictiveness across the three models is unchanged. The model class is defined as in Section 3.3. The structural mapping $S$ corresponds to the mixed logit functional form in the BLP framework, while for the multinomial logit and nested logit specifications we replace $S$ with the corresponding functional forms used in the previous section. See Appendix F for details.
@@
To make the results comparable to the no-endogeneity case in Section 5.2, we use the same definition for the eligible sets–“NP Both,” “NP Mean,” and “NP Individual,” which differ in whether the mean utility and/or individual heterogeneity are allowed to be nonparametric–along with its corresponding evaluation distribution, to generate the pseudo-true prediction rules. We draw 50 functions from the evaluation distribution $\lambda_{\mathcal{F}}.$
@@
where $\bar{h}(Z)=\mathbb{E}[h(X,Z)|Z]$. In the implementation, $\bar{h}$ is obtained by projecting $h$ onto higher-order polynomial functions of the instruments $Z$. When computing the infimum over $h$, we draw 20 candidate $h$-functions from a Gaussian process prior with Matérn-3/2 kernel.  [^9] Robustness checks that vary the kernel choice and the function-draw procedure yield similar results; see Online Appendix F.3.
Results.
Table 4 reports restrictiveness for the discrete-choice models when price is treated as endogenous. Relative to the no-endogeneity case, restrictiveness is substantially higher for each model under each eligible set. For example, under eligible set “NP Both”, restrictiveness for MNL increases from 0.154 to 0.758, and for MXL from 0.112 to 0.679. This reflects the additional constraints imposed by the moment conditions: each structural model must now match both its functional form and the exogeneity restrictions, which makes it harder to approximate the pseudo-true share.
Across all three eligible sets, MXL is the least restrictive model, while MNL and NL yield nearly identical restrictiveness. This contrasts with the no-endogeneity case, where both NL and MXL are less restrictiveness than MNL; here, the nesting parameter in NL does not play a role in relaxing restrictiveness once endogeneity restrictions are imposed. Moreover, as in the no-endogeneity case, the role of the mean utility component remains central. Restrictiveness under the “NP Mean” eligible set is close to that under “NP Both” for all models. By contrast, the “NP Individual” eligible set delivers somewhat lower restrictiveness, but the levels remain high, ranging from 0.587 to 0.631. Moment restrictions prevent a parametric mean utility component from yielding restrictiveness values close to zero.
Table 4: Restrictiveness of Multinomial choice Models (With Endogeneity)
Eligible Set Model Restr. SE
NP Both MNL 0.758 0.012
NP Both NL 0.758 0.012
NP Both MXL 0.679 0.020
NP Individual MNL 0.631 0.009
NP Individual NL 0.631 0.009
NP Individual MXL 0.587 0.014
NP Mean MNL 0.729 0.012
NP Mean NL 0.729 0.012
NP Mean MXL 0.631 0.011
Table 5 reports completeness for the three models. Mixed logit achieves slightly higher completeness than MNL and NL, while MNL and NL remain nearly identical.
Table 5: Completeness of Multinomial Choice Models (With Endogeneity)
Model Complete. SE
MNL 0.301 0.035
NL 0.302 0.035
MXL 0.335 0.040
@@ 6 Conclusion
In this paper, we further develop the notion of restrictiveness proposed in Fudenberg, Gao and Liang (2026) to functional and structural settings, which substantially enlarges the scope under which the restrictiveness measure can be used to evaluate the a priori economic structures imposed by a specific economic or econometric model. Restrictiveness, together with completeness (Fudenberg, Kleinberg, Liang and Mullainathan, 2022), provide a general-purpose toolkit for theoretical and applied researchers to evaluate the “value added” by economic theory, or more generally any restrictions imposed by any domain-specific structures, beyond purely statistical aspects of their models.
Our work also suggests several directions of future research: (1) Investigate the restrictiveness-completeness frontier in a larger variety of economic applications, and provide a more comprehensive view and understanding of the “value added” by economic theory in various empirical settings. (2) Investigate how to incorporate recent advances in the Bayesian nonparametrics and machine learning literature to make the procedure of sampling from the prior distribution $\lambda_{\mathcal{F}}$ more computationally efficient and robust, especially under various shape restrictions. (3) Explore the use of restrictiveness as a regularization device in machine learning.
We think that direction (3) is particularly interesting. The idea is to extend our framework by regularizing empirical risk minimization with a penalty term (1r)(1-r), where $r$ measures the model’s restrictiveness. Unlike syntactic complexity measures (such as parameter counting in AIC or BIC), such a penalty would favor models based on their population-level structural content and adherence to economic constraints. This approach could be applied both to selection across non-nested model classes and to regularization within parametric families, favoring regions of the parameter space that induce more restrictive prediction rules. An important empirical question is whether such restrictiveness-based regularization lead to better finite-sample performance in estimation and prediction than standard complexity penalties.
@@ Appendix A Proof of Proposition 1
$\mathbb{E}\left[(\overline{m}_{\theta}(X)-\overline{g}(X))(u(Y_{c},X)-v(Y_{c},X))\right]=\mathbb{E}_{X}\left[(\overline{f}_{\theta}-\overline{g})\underbrace{\mathbb{E}[u-v|X]}_{=0}\right]=0.$
@@
$\inf_{f\in\mathcal{F}_{\Theta}}d(f,g)$ $=\inf_{\theta\in\Theta}\mathbb{E}_{P_{X}}\left[(\overline{f}_{\theta}(X)-\overline{g}(X))^{2}\right]+\inf_{h\in\mathcal{F}}\mathbb{E}_{P_{X,X}}\left[(u(Y_{c},X)-v(Y_{c},X))^{2}\right]$
@@
Appendix B Structural-Form Restrictiveness
Alternatively, we may want to focus on the structural form as a mapping from endogenous and exogenous covariates to outcomes instead of the reduced form. In addition, there are many partially specified structural models (often identified via the use of IV exogeneity conditions) that do not admit a reduced-form representation. In such cases, we a notion of structural-form restrictiveness under the following assumption.
@@ Appendix B Structural-Form Restrictiveness
for some known mapping $f_{\theta}$ and parameter $\theta_{0}\in\Theta$, with $Y_{o,i}$ denoting the outcome variable and $Y_{c.i}$ denoting the endogenous covariates.
@@
The reduced-form restrictiveness $r_{\mathrm{RF}}$ of a model is therefore a property of the image of a structural form model (2) parametrized by $\theta\in\Theta$ under the equilibrium operator $T$, while the structural-form restrictiveness $r_{\mathrm{SF}}$ is a property of the structural class structural form model (2) itself, before applying the equilibrium operator $T$. In general, these two need not coincide.
@@
Appendix C Inference with Estimated Discrepancies
As in Ellis and Neff (2025), we assume that for $\lambda_{\mathcal{F}}$-almost every $f\in F$ the discrepancy admits a moment representation $d(f_{\theta},f)\;=\;\mathbb{E}_{P_{X}}\left[g(X,\theta;f)\right],$ for a measurable function $g:{\bf X}\times\Theta\times\mathcal{F}\to\mathbb{R}$, and similarly $d(f_{\mathrm{base}},f)\;=\;\mathbb{E}_{P_{X}}\left[g_{\mathrm{base}}(X;f)\right]$ for some measurable $g_{\mathrm{base}}:{\bf X}\times\mathcal{F}\to\mathbb{R}$.
Assumption 6 (Finite Moments).
Given $S_{n}$ we define the empirical analogs
and the profiled empirical discrepancy
Assumption 7 (Asymptotic Linearity).
and the combined influence function
Theorem 1 (Asymptotic Normality of $\widehat{r}_{n,M}$).
Under Assumption 7. Fix $M\geq 1$ and independent draws $f_{1},\dots,f_{M}\sim\lambda_{\mathcal{F}}$, and define $\mu_{1,M}$, $\mu_{0,M}$ and $r_{M}$ as above. Then, conditional on $(f_{1},\dots,f_{M})$,
as $n\to\infty$, where $\psi_{M}$ is given by (28).
Proof.
By Assumption 7,
$\widehat{\mu}_{1,n,M}$ $=\frac{1}{M}\sum_{m=1}^{M}d_{n}(F_{\Theta},f_{m})$
$=\mu_{1,M}+\frac{1}{n}\sum_{i=1}^{n}\Phi_{i}^{(1)}+o_{p}(n^{-1/2}),$
and similarly
Therefore
By the delta method,
Define the empirical analogs
and
The plug-in variance estimator is
Appendix D Rademacher Complexity and VC Dimension
@@
Appendix E Limit of Gaussian-Process Learning Curve
The analysis of the limit of the learning curve in Proposition 2 is related to the literature on learning curves for Gaussian process (GP) regression. To articulate the connection and the difference, it is useful to separate the roles of approximation, estimation, and irreducible noise. In GP regression, one typically takes squared loss $\ell(y,y^{\prime})=(y-y^{\prime})^{2}$, so $d(f,g)$ is an $L^{2}(P_{X})$ prediction error. The average generalization error of the posterior mean (or BLUP) can then be written in exactly the form of our learning curve $L_{n}(F_{\Theta},A)$, with the average taken over both training samples and draws of the pseudo-truth function $f$ from a GP prior.
Le Gratiet and Garnier (2015) consider this setting with $X_{i}\sim\mu$ on $\mathbb{R}^{d}$, a zero-mean GP prior with covariance kernel $k$, and noisy observations $Z(x_{i})+\varepsilon_{i}$ where the noise variance scales as $\mathrm{Var}(\varepsilon_{i})=n\tau$. They study the integrated mean squared error (IMSE) of the BLUP,
@@ Appendix E Limit of Gaussian-Process Learning Curve
$\mathrm{IMSE}_{n}\;\xrightarrow{p}\;\sum_{p\geq 0}\frac{\tau\lambda_{p}}{\tau+\lambda_{p}},\text{ as }n\to\infty.$ (31)
@@
In our notation, this corresponds to the following specialization. Let $\mathcal{F}$ be the set of sample paths of the GP, let $\lambda_{\mathcal{F}}$ be the GP prior, let $P_{X}=\mu$, and take $\ell(y,y^{\prime})=(y-y^{\prime})^{2}$, so that $d(f,g)$ is squared $L^{2}(\mu)$ distance between functions. Let $\mathcal{A}$ be the BLUP estimator based on noisy observations. Then $L_{n}(\mathcal{F}_{\Theta},\mathcal{A})$ is exactly the GP learning curve (IMSE) when we average over both $f\sim\lambda_{\mathcal{F}}$ and samples. By Proposition 2, in a noise-free ($\tau=0$) and risk-consistent setting, $\lim_{n}L_{n}(\mathcal{F}_{\Theta},A)$ must equal $\mathbb{E}_{f\sim\lambda_{\mathcal{F}}}[d(\mathcal{F}_{\Theta},f)]$, i.e. the average approximation error of $\mathcal{F}_{\Theta}$ relative to $(\mathcal{F},\lambda_{\mathcal{F}},d)$.
@@
Appendix F Details and Additional Results for Section 5
F.1 Details for Section 5.1
@@ Appendix F Details and Additional Results for Section 5
Table A.2 reports restrictiveness under this alternative kernel. Relative to the baseline results in Table 2, restrictiveness decreases for all three models. Importantly, the qualitative ranking of models remains unchanged: MNL is the most restrictive model, while NL and MNL have similar restrictiveness.
@@
NP Both MNL 0.132 0.009
NP Both NL 0.089 0.005
NP Both MXL 0.089 0.005
NP Individual MNL 0.000 0.000
NP Individual NL 0.000 0.000
NP Individual MXL 0.001 0.000
NP Mean MNL 0.133 0.009
NP Mean NL 0.091 0.005
NP Mean MXL 0.095 0.005
@@
As a robustness check, we replace the GP draws for both the common component $f$ and the individual component $f_{i}$ with cubic spline basis draws. Specifically, in the sampling procedure outlined in F.2, the latent process $h$ is now generated from a spline-basis draw. We construct a truncated-power cubic basis
@@
NP Both MNL 0.251 0.009
NP Both NL 0.211 0.006
NP Both MXL 0.211 0.007
NP Individual MNL 0.002 0.000
NP Individual NL 0.002 0.000
NP Individual MXL 0.002 0.000
NP Mean MNL 0.244 0.008
NP Mean NL 0.205 0.006
NP Mean MXL 0.206 0.006
F.3 Details for Section 5.3
@@
As a robustness check, we report restrictiveness results using three instruments that are most strongly correlated with the endogenous regressor in Table A.4. Relative to the specification with fewer instruments (Table 4), restrictiveness increases across all models. This is consistent with theory, as adding an additional moment condition further constrains the model’s ability to fit the pseudo data. Importantly, the qualitative ranking of models remains unchanged: MNL and NL exhibit nearly identical restrictiveness, while MXL remains less restrictive than the other two.
@@
NP Both MNL 0.789 0.012
NP Both NL 0.789 0.012
NP Both MXL 0.682 0.021
NP Individual MNL 0.670 0.008
NP Individual NL 0.670 0.008
NP Individual MXL 0.605 0.014
NP Mean MNL 0.799 0.010
NP Mean NL 0.799 0.010
NP Mean MXL 0.663 0.010
@@
NP Both MNL 0.754 0.013
NP Both NL 0.754 0.013
NP Both MXL 0.672 0.021
NP Individual MNL 0.636 0.009
NP Individual NL 0.636 0.009
NP Individual MXL 0.588 0.014
NP Mean MNL 0.730 0.013
NP Mean NL 0.730 0.013
NP Mean MXL 0.634 0.010
@@
NP Both MNL 0.745 0.012
NP Both NL 0.745 0.012
NP Both MXL 0.652 0.010
NP Individual MNL 0.462 0.006
NP Individual NL 0.462 0.006
NP Individual MXL 0.442 0.005
NP Mean MNL 0.746 0.012
NP Mean NL 0.746 0.012
NP Mean MXL 0.655 0.011

### v3

@@ 5.3 Multinomial Choice with Endogeneity
We now turn to the empirically important case where product characteristics, particularly price, are endogenous. We address this endogeneity using the BLP instruments constructed in Nevo (2000), selecting the two instruments that are most strongly correlated with price from the full set of 20.  [^8] In Online Appendix S.4.3, we also consider selecting the three instruments most strongly correlated with price; the qualitative ranking of restrictiveness across the three models is unchanged. The model class is defined as in Section 3.3. The structural mapping $S$ corresponds to the mixed logit functional form in the BLP framework, while for the multinomial logit and nested logit specifications we replace $S$ with the corresponding functional forms used in the previous section. See Online Appendix S.4.3 for details.
@@
To make the results comparable to the no-endogeneity case in Section 5.2, we use the same definition for the eligible sets–“NP Both,” “NP Mean,” and “NP Individual,” which differ in whether the mean utility and/or individual heterogeneity are allowed to be nonparametric–along with its corresponding evaluation distribution, to generate the pseudo-true prediction rules. We draw 500 functions from the evaluation distribution $\lambda_{\mathcal{F}}.$
@@
where $\bar{h}(Z)=\mathbb{E}[h(X,Z)|Z]$. In the implementation, $\bar{h}$ is obtained by projecting $h$ onto higher-order polynomial functions of the instruments $Z$. When computing the infimum over $h$, we draw 100 candidate $h$-functions from a Gaussian process prior with Matérn-3/2 kernel.  [^9] Robustness checks that vary the kernel choice and the function-draw procedure yield similar results; see Online Appendix S.4.3.
Results
In Panel A of Table 2, the “With Endogeneity” columns report restrictiveness for the discrete-choice models when price is treated as endogenous. Relative to the no-endogeneity case, restrictiveness is substantially higher for each model under each eligible set, as recorded in the “Change” column. For example, under eligible set “NP Both”, restrictiveness for MNL increases from 0.154 to 0.745, and for MXL from 0.112 to 0.667. This reflects the additional constraints imposed by the moment conditions: each structural model must now match both its functional form and the exogeneity restrictions, which makes it harder to approximate the pseudo-true share.
Across all three eligible sets, MXL is the least restrictive model, while MNL and NL yield nearly identical restrictiveness. This contrasts with the no-endogeneity case, where both NL and MXL are less restrictive than MNL; here, the nesting parameter in NL does not play a role in relaxing restrictiveness once endogeneity restrictions are imposed. Moreover, as in the no-endogeneity case, the role of the mean utility component remains central. Restrictiveness under the “NP Mean” eligible set is close to that under “NP Both” for all models. By contrast, the “NP Individual” eligible set delivers somewhat lower restrictiveness, but the levels remain high, ranging from 0.548 to 0.594. Moment restrictions prevent a parametric mean utility component from yielding restrictiveness values close to zero.
In Panel B of Table 2, the “With Endogeneity” columns report completeness for the three models. Mixed logit achieves slightly higher completeness than MNL and NL, while MNL and NL remain nearly identical.
@@ 6 Conclusion
This paper establishes that model restrictiveness can be defined, computed, and interpreted in functional and structural settings far beyond those considered in prior work. Doing so yields economically meaningful findings that finite-support evaluations miss: most strikingly, IV moment conditions can raise restrictiveness more dramatically than functional form restrictions alone, as when mixed logit restrictiveness increases from 0.11 to 0.67 in our BLP application.
Together with the completeness measure of Fudenberg, Kleinberg, Liang and Mullainathan (2022), restrictiveness gives researchers a two-dimensional lens for model evaluation: how much a model rules out, and how much of what matters it captures. Neither dimension alone is sufficient: a model can be highly restrictive yet empirically uninformative, or highly complete yet theoretically vacuous. The restrictiveness–completeness frontier, first proposed in Fudenberg, Gao and Liang (2026), thus provides a principled basis for comparing models that would be difficult to rank on either criterion alone.
Our work suggests three directions for future research. First, mapping the restrictiveness–completeness frontier across a wider range of structural models, including dynamic models and models with strategic interaction, would provide a more comprehensive picture of where economic theory adds value beyond statistical fit. Second, more computationally efficient sampling algorithms for constrained functional spaces, drawing on recent advances in Bayesian nonparametrics, would make the framework more accessible to applied researchers working with high-dimensional covariates. Third, and most intriguing, is the possibility of using restrictiveness as a regularization device. Standard complexity penalties such as AIC and BIC count parameters, a syntactic measure that ignores whether those parameters encode economically meaningful structure. A restrictiveness-based penalty that favors models with high population-level structural content would instead reward models that rule out economically implausible patterns, regardless of their parameter count. Such a penalty could be applied both to selection across non-nested model classes and to regularization within parametric families, and whether it improves finite-sample estimation and prediction relative to standard penalties is an important open question.
@@ Appendix A Proof of Proposition 1
$\mathbb{E}\left[(\overline{f}_{\theta}(X)-\overline{g}(X))(u(Y_{c},X)-v(Y_{c},X))\right]=\mathbb{E}_{X}\left[(\overline{f}_{\theta}-\overline{g})\underbrace{\mathbb{E}[u-v|X]}_{=0}\right]=0.$
@@
$\inf_{f\in\mathcal{F}_{\Theta}}d(f,g)$ $=\inf_{\theta\in\Theta}\mathbb{E}_{P_{X}}\left[(\overline{f}_{\theta}(X)-\overline{g}(X))^{2}\right]+\inf_{h\in\mathcal{F}}\mathbb{E}_{P_{X}}\left[(u(Y_{c},X)-v(Y_{c},X))^{2}\right]$
@@
Appendix B Inference with Estimated Discrepancies
As in Ellis and Neff (2025), we assume that for $\lambda_{\mathcal{F}}$-almost every $f\in\mathcal{F}$ the discrepancy admits a moment representation $d(f_{\theta},f)\;=\;\mathbb{E}_{P_{X}}\left[g(X,\theta;f)\right],$ for a measurable function $g:{\bf X}\times\Theta\times\mathcal{F}\to\mathbb{R}$, and similarly $d(f_{\mathrm{base}},f)\;=\;\mathbb{E}_{P_{X}}\left[g_{\mathrm{base}}(X;f)\right]$ for some measurable $g_{\mathrm{base}}:{\bf X}\times\mathcal{F}\to\mathbb{R}$.
Assumption 5 (Finite Moments).
Given $S_{n}$ we define the empirical analogs
and the profiled empirical discrepancy
Assumption 6 (Asymptotic Linearity).
and the combined influence function
Theorem 1 (Asymptotic Normality of $\widehat{r}_{n,M}$).
Under Assumption 6, fix $M\geq 1$ and independent draws $f_{1},\dots,f_{M}\sim\lambda_{\mathcal{F}}$, and define $\mu_{1,M}$, $\mu_{0,M}$ and $r_{M}$ as above. Then, conditional on $(f_{1},\dots,f_{M})$,
as $n\to\infty$, where $\psi_{M}$ is given by (25).
Proof.
By Assumption 6,
$\widehat{\mu}_{1,n,M}$ $=\frac{1}{M}\sum_{m=1}^{M}d_{n}(F_{\Theta},f_{m})$
$=\mu_{1,M}+\frac{1}{n}\sum_{i=1}^{n}\Phi_{i}^{(1)}+o_{p}(n^{-1/2}),$
and similarly
Therefore
By the delta method,
Define the empirical analogs
and
The plug-in variance estimator is
Online Appendix
Appendix S.1 Structural-Form Restrictiveness
Alternatively, we may want to focus on the structural form as a mapping from endogenous and exogenous covariates to outcomes instead of the reduced form. In addition, there are many partially specified structural models (often identified via the use of IV exogeneity conditions) that do not admit a reduced-form representation. In such cases, we define a notion of structural-form restrictiveness under the following assumption.
@@ Appendix B Inference with Estimated Discrepancies
for some known mapping $f_{\theta}$ and parameter $\theta_{0}\in\Theta$, with $Y_{o,i}$ denoting the outcome variable and $Y_{c,i}$ denoting the endogenous covariates.
@@
The reduced-form restrictiveness $r_{\mathrm{RF}}$ of a model is therefore a property of the image of a structural-form model (2) parametrized by $\theta\in\Theta$ under the equilibrium operator $T$, while the structural-form restrictiveness $r_{\mathrm{SF}}$ is a property of the structural class structural-form model (2) itself, before applying the equilibrium operator $T$. In general, these two need not coincide.
@@
Appendix S.2 Rademacher Complexity and VC Dimension
@@
Appendix S.3 Gaussian-Process Learning Curve: Details
This section provides further details on the connection between restrictiveness and the Gaussian process (GP) learning curve, complementing the discussion in Section 4.3.
It is useful to separate the roles of approximation, estimation, and irreducible noise. In GP regression, one typically takes squared loss $\ell(y,y^{\prime})=(y-y^{\prime})^{2}$, so $d(f,g)$ is an $L^{2}(P_{X})$ prediction error. Le Gratiet and Garnier (2015) consider the setting with $X_{i}\sim\mu$ on $\mathbb{R}^{d}$, a zero-mean GP prior with covariance kernel $k$, and noisy observations $Z(x_{i})+\varepsilon_{i}$ where the noise variance scales as $\mathrm{Var}(\varepsilon_{i})=n\tau$. They study the integrated mean squared error (IMSE) of the best linear unbiased predictor (BLUP),
@@
$\mathrm{IMSE}_{n}\;\xrightarrow{p}\;\sum_{p\geq 0}\frac{\tau\lambda_{p}}{\tau+\lambda_{p}},\text{ as }n\to\infty,$ (31)
@@
In our notation, this corresponds to the following specialization. Let $\mathcal{F}$ be the set of sample paths of the GP, let $\lambda_{\mathcal{F}}$ be the GP prior, let $P_{X}=\mu$, and take $\ell(y,y^{\prime})=(y-y^{\prime})^{2}$, so that $d(f,g)$ is the squared $L^{2}(\mu)$ distance between functions. Let $\mathcal{A}$ be the BLUP estimator based on noisy observations. Then $L_{n}(\mathcal{F}_{\Theta},\mathcal{A})$ is exactly the GP learning curve (IMSE) when we average over both $f\sim\lambda_{\mathcal{F}}$ and samples. By Proposition 2, in a noise-free ($\tau=0$) and risk-consistent setting, $\lim_{n}L_{n}(\mathcal{F}_{\Theta},\mathcal{A})$ must equal $\mathbb{E}_{f\sim\lambda_{\mathcal{F}}}[d(\mathcal{F}_{\Theta},f)]$, i.e. the average approximation error of $\mathcal{F}_{\Theta}$ relative to $(\mathcal{F},\lambda_{\mathcal{F}},d)$.
@@
Appendix S.4 Details and Additional Results for Section 5
S.4.1 Details for Section 5.1
@@
Table A.2 reports restrictiveness under this alternative kernel. Relative to the baseline results in Panel A of Table 2, restrictiveness decreases for all three models. Importantly, the qualitative ranking of models remains unchanged: MNL is the most restrictive model, while NL and MNL have similar restrictiveness.
@@
NP Both MNL 0.132 (0.009)
NP Both NL 0.089 (0.005)
NP Both MXL 0.088 (0.005)
NP Individual MNL 0.000 (0.000)
NP Individual NL 0.000 (0.000)
NP Individual MXL 0.000 (0.000)
NP Mean MNL 0.132 (0.009)
NP Mean NL 0.090 (0.005)
NP Mean MXL 0.091 (0.005)
@@
As a robustness check, we replace the GP draws for both the common component $f$ and the individual component $f_{i}$ with cubic spline basis draws. Specifically, in the sampling procedure outlined in S.4.2, the latent process $h$ is now generated from a spline-basis draw. We construct a truncated-power cubic basis
@@
NP Both MNL 0.253 (0.009)
NP Both NL 0.213 (0.006)
NP Both MXL 0.213 (0.007)
NP Individual MNL 0.002 (0.000)
NP Individual NL 0.002 (0.000)
NP Individual MXL 0.001 (0.000)
NP Mean MNL 0.241 (0.008)
NP Mean NL 0.203 (0.005)
NP Mean MXL 0.203 (0.005)
S.4.3 Details for Section 5.3
@@
As a robustness check, we report restrictiveness results using three instruments that are most strongly correlated with the endogenous regressor in Table A.4. Relative to the specification with fewer instruments (Panel A of Table 2), restrictiveness increases across all models. This is consistent with theory, as adding an additional moment condition further constrains the model’s ability to fit the pseudo data. Importantly, the qualitative ranking of models remains unchanged: MNL and NL exhibit nearly identical restrictiveness, while MXL remains less restrictive than the other two.
@@
NP Both MNL 0.779 (0.013)
NP Both NL 0.779 (0.013)
NP Both MXL 0.673 (0.022)
NP Individual MNL 0.651 (0.007)
NP Individual NL 0.651 (0.007)
NP Individual MXL 0.589 (0.014)
NP Mean MNL 0.800 (0.010)
NP Mean NL 0.800 (0.010)
NP Mean MXL 0.665 (0.010)
@@
NP Both MNL 0.742 (0.014)
NP Both NL 0.742 (0.014)
NP Both MXL 0.663 (0.022)
NP Individual MNL 0.596 (0.009)
NP Individual NL 0.596 (0.009)
NP Individual MXL 0.544 (0.015)
NP Mean MNL 0.698 (0.009)
NP Mean NL 0.698 (0.009)
NP Mean MXL 0.607 (0.008)
@@
NP Both MNL 0.714 (0.009)
NP Both NL 0.714 (0.009)
NP Both MXL 0.626 (0.008)
NP Individual MNL 0.415 (0.005)
NP Individual NL 0.415 (0.005)
NP Individual MXL 0.398 (0.004)
NP Mean MNL 0.716 (0.010)
NP Mean NL 0.716 (0.010)
NP Mean MXL 0.628 (0.008)